OpenStudy (d_elam):

Rewrite in terms of sin(x) and cos(x). sin(x+7π/4)

satellite73 (satellite73):

\[A\cos(x)+B\sin(x)=\sqrt{A^2+B^2}\sin(x+\theta)\] you have to go from the right to the left, which is kind of odd but doable

OpenStudy (d_elam):

So I plug what I know into the right side of that equation?

satellite73 (satellite73):

you need A and B

satellite73 (satellite73):

\[\sqrt{A^2+B^2}=1\] in your example

OpenStudy (sshayer):

\[\sin \left( x+\frac{ 7 \pi }{ 4 } \right)=\sin x \cos \frac{ 7 \pi }{ 4 }+\cos x \sin \frac{ 7 \pi }{ 4 }\] \[=\sin x \cos \left( 2 \pi-\frac{ \pi }{ 4 } \right)+\cos x \sin \left( 2 \pi-\frac{ \pi }{ 4 } \right)\] \[=\sin x \cos \left( -\frac{ \pi }{ 4 } \right)+\cos x \sin \left( -\frac{ \pi }{ 4 } \right)\] \[=\sin x \cos \frac{ \pi }{ 4 }-\cos x \sin \frac{ \pi }{ 4 }\] \[=\frac{ 1 }{ \sqrt{2} }\sin x-\frac{ 1 }{ \sqrt{2} }\cos x \]